### v2

# Need for zkSpeed: Accelerating HyperPlonk for Zero-Knowledge Proofs

Abstract.
Zero-Knowledge Proofs (ZKPs) are a rapidly growing technique for privacy-preserving and verifiable computation. ZKPs enable one party (a prover: ) to prove to another (a verifier: ) that a statement is true or correct without revealing any additional information. This powerful capability has led to ZKPs being applied and proposed for application in blockchain technologies, verifiable machine learning, and electronic voting. However, ZKPs have yet to see widespread, ubiquitous adoption due to the exceptionally high computational complexity of the proving process. Naturally, there has been recent work to accelerate ZKP primitives and protocols using GPUs and ASICs. However, the protocols considered so far face one of two challenges: they require a trusted setup for each new application or generate large proofs with high verification costs, limiting their applicability in scenarios with numerous verifiers or strict verification time constraints. HyperPlonk is a state-of-the-art ZKP protocol that supports both one-time, universal setup and small proof sizes/verification costs expected by publicly verifiable, consensus-based systems (e.g., blockchain). While HyperPlonk’s setup and verifier properties are highly desirable, the proving phase is costly. A HyperPlonk prover must compute on large bitwidths (e.g., 255-381b) and polynomials (e.g., of degree 224), employs computationally (e.g., MSM) and bandwidth (e.g., SumCheck) intensive kernels, and the complete protocol comprises many steps, each constituting distinct kernels. We present an accelerator, zkSpeed, to address these challenges and effectively accelerate HyperPlonk. zkSpeed provides hardware support for all major primitives (e.g., SumCheck and Multi-Scalar Multiplications (MSMs)) and judiciously schedules each protocol phase onto the allocated hardware. We leverage high-level synthesis to thoroughly explore and optimize the hardware design tradeoffs of each unit. These are then input into a full-chip simulator for large-scale design space exploration to optimize all aspects of the architecture in unison. Our Pareto analysis demonstrates that with a 366mm2 chip and 2 TB/s of off-chip bandwidth, zkSpeed is able to accelerate the entire proof generation by 801 (geomean) over a CPU baseline.
1. Introduction
Zero-Knowledge Proofs (ZKPs) are a privacy-preserving computation technique that enable one party (a prover ) to prove to another (a verifier ) that a computation (possibly with secret inputs) was performed correctly. The proof itself reveals no information about the (secret) inputs nor the function, and it can be verified much faster than the compute time of the original computation. Proof systems can be characterized by several important criteria, including: prover time, verifier time, proof size, protocol setup, and cryptographic assumptions required for the proof to be secure. Each deployment of ZKPs can result in a different set of criteria requirements. For example, private transactions (Ben-Sasson et al., 2014) require one proof per transaction be posted on a blockchain and distributed to all blockchain nodes, thus prioritizing small proof size.
Recent years have seen significant advances and heterogeneity in proof systems (Groth, 2016; Gabizon et al., 2019; Chen et al., 2022; Xie et al., 2022; Setty, 2020; Abbaszadeh et al., 2024). Each optimizes for a distinct set of criteria constraints, and we note that in all systems, prover time is vital to optimize for, as faster provers enable new applications and proving is much slower than verification (lib, 2018; Chen et al., 2022). To understand the ZKP space, consider Groth16 (Groth, 2016) and Orion (Xie et al., 2022), two ZKPs with distinct characteristics that have also recently been accelerated (Zhang et al., 2021; Daftardar et al., 2024; Samardzic et al., 2024). Groth16 produces very short proofs (192 bytes) and enables millisecond verification latency, independent of proved computation size. However, it relies on a strong cryptographic assumption: a circuit-specific trusted setup where a trusted party generates keys using secret randomness. If this randomness is not properly discarded or the party is malicious, the system’s security is compromised, allowing false statements to be proven. Hence, ZCash (Ben-Sasson et al., 2014) and other applications have moved away from circuit-specific trusted-setup protocols (Nelson, 2022). On the other hand, Orion does not require a trusted setup or rely on computationally expensive elliptic curve cryptography, enabling a faster prover—but at the cost of large (8 MB) proofs. To put this in perspective, Orion’s proofs are roughly 4 larger than the maximum block size observed in Ethereum cryptocurrency (Monnot, 2024). Thus, a private transaction with an Orion proof would not fit into a single Ethereum block.
HyperPlonk (Chen et al., 2022) is a recent system gaining attention for the ZKP criteria it optimizes for, which are desirable in many applications. It has small proofs (typically 5 KB), low verification complexity, and supports universal trusted setup (Groth et al., 2018), which runs once for all time and then can be reused by new applications. A major contribution of HyperPlonk is its elimination of the Number Theoretic Transform (NTT), which is used by Groth16 and others, and instead its usage of SumCheck, which lowers the asymptotic runtime from (NTT) to (SumCheck). Here, represents the size of the underlying computation, and thus a lower runtime implies larger problems may be applied to ZKPs; something we cannot do today due to long runtime. Despite the improved runtime, the proving phase of HyperPlonk is still slow, running on the order of minutes to hours for some applications.
Accelerating HyperPlonk presents many challenges. First, the prover works over large, computationally demanding data structures, with polynomials on the order of 217-224 and bitwidths ranging from 255-381. Second, the protocol constitutes four major phases of computation, each comprising a different set of kernel functions. Kernels range from small and large compute-bound problems (e.g., SHA3 and MSM, respectively) to memory-bound (e.g., SumCheck). Compared to more traditionally accelerated kernels (e.g., NTT, matrix multiply, and convolution), HyperPlonk’s kernels are also less well understood. Finally, there is heterogeneity in data access patterns across kernels and phases. Some data structures are reused, either between or within kernels, and some are not. This raises questions on how to connect and schedule kernels as well as how to stage data to overlap communication with computation. The complexity of HyperPlonk requires careful design, new hardware optimizations, and thorough design space exploration to identify architectures with area-performance tradeoffs that both meet the computational needs of applications and justify the costs of hardware acceleration.
In this paper we present zkSpeed, a modular HyperPlonk accelerator that overcomes the challenges outlined above. zkSpeed comprises eight unique kernel accelerators, called units. Each unit supports required operators (e.g., large-word modular arithmetic) natively in hardware over a programmable problem size. The units collectively implement the complete zkSpeed chip architecture, which includes shared and unit-local scratchpads to capture reuse, capable of processing entire HyperPlonk proofs. Units and scratchpads communicate via a multi-channel shared bus, and we propose a schedule of each (sequential) HyperPlonk protocol phase onto zkSpeed accordingly. The schedule balances off-chip bandwidth constraints with on-chip reuse from scratchpads, unit computational throughput, and bus resource contention with a goal of overlapping off-chip communication with computation whenever possible. All unit designs and bandwidth constraints are configurable and can be tailored to use the least amount of area needed to achieve this goal. To optimize all units under a given off-chip bandwidth, we developed a zkSpeed simulator capable of sweeping over a wide range of design parameter settings. We consider thousands of unique zkSpeed designs. Analyzing the resulting Pareto frontier, we conclude that a 366mm2 chip with 2 TB/s of off-chip bandwidth can speedup proof generation by 801 (geomean) over a CPU baseline.
This paper makes the following contributions:
A high-throughput, fully-pipelined accelerator to handle three flavors of SumCheck for HyperPlonk-based proofs.
A novel implementation for modular inversion to compute fraction polynomials not seen in other ZKP protocols.
Additional optimizations–including resource sharing of modular multipliers across units and on-chip memory compression schemes–that save upwards of 50% area per unit and up to 85% in bandwidth utilization.
A comprehensive design space exploration of all hardware units to investigate design tradeoffs as we scale to high-performance designs and advanced memory technologies.
A full-chip design that achieves 801 gmean speedup over CPU at iso-compute area.
2. Background
In this section, we describe state-of-the-art ZKPs and introduce the mathematical principles underlying the key kernels of HyperPlonk.
2.1. zkSNARKs
Today, the state-of-the-art ZKP protocols are zero-knowledge Succinct Non-interactive Arguments of Knowledge (zkSNARKs). zkSNARKs have three properties: (i) zero-knowledge, i.e., the proof does not reveal any information about the secret witness ; (ii) succinct, i.e., the proof has a few hundreds of bytes; and (iii) non-interactive, i.e., sends the proof to in one exchange.
zkSNARKs like HyperPlonk use polynomials to encode the correct execution of the target program. Polynomials can encode complex computations and constraints in a compact form, reducing the computational burden on to a few checks, instead of requiring inspection of individual operations in the program. In HyperPlonk, the two most time-consuming kernels are SumCheck and Multi-Scalar Multiplication (MSM). The SumCheck kernel lets demonstrate knowledge of a polynomial by verifying that its properties hold over a large set of inputs without revealing, e.g., the polynomial’s coefficients. MSMs ensure that is bound to that polynomial, preventing manipulation of the polynomial in the proof.
2.2. The SumCheck Kernel
SumCheck (Thaler, 2022) is an interactive protocol between a prover and a verifier . demonstrates to that it has correctly computed the sum of a polynomial over the boolean hypercube, i.e., over all Boolean (0/1) assignments of its variables.
Given a multivariate polynomial , where 111each variable can be an integer modulo a prime number ., wants to prove to that it correctly computed the sum
and engage in a multi-round protocol. In Round 1, computes , a univariate polynomial of by summing over all binary values of the remaining variables:
then sends coefficients or evaluations of to . checks that , and if so, generates a random challenge and sends it to .
In Round 2, asks to prove that
which is simply an instance of SumCheck, except over a ()-variate polynomial. Thus, in Round 2, computes and returns
to and the protocol repeats recursively for a total of rounds. If all checks pass, accepts ’s claim about .
2.3. Multilinear Polynomials
ZKPs like HyperPlonk use multilinear polynomials, that are linear in each of their variables. For example,
is multilinear because the maximum degree of individual variables is one. ZKP protocols use a general representation of multilinear polynomials, shown below for a 2-variable multilinear polynomial:
This representation defines polynomials using evaluations at specific input points (e.g., ) rather than using coefficients of symbolic terms (e.g., ). Just as a degree- polynomial is uniquely determined by coefficients, it can also be uniquely defined by its values at distinct points. This representation then maps cleanly to hardware in the form of lookup tables. For example, for a 2-variable multilinear polynomial, the binary values of index into the table to get the corresponding evaluation. In general, a multilinear polynomial with variables can be stored in a lookup table of entries. As we will discuss in Section 3, HyperPlonk uses multilinear polynomials as building blocks for higher-degree polynomials on which we then perform SumCheck and other computations. In the rest of the paper, we will use the term MLE table to refer to these lookup data structures (MLE stands for “multilinear extensions” (Thaler, 2022); the word “extension” indicates that these polynomials can be evaluated at non-boolean, or extended, values).
2.4. The MSM Kernel
MSMs are dot products between a vector of scalars and a vector of 2D or 3D points on an elliptical curve, e.g., . MSMs are used in ZKPs to perform commitments. A commitment is a cryptographic primitive that binds a prover to a value without revealing it. In zkSNARK protocols like HyperPlonk (Chen et al., 2022), scalars are the polynomial evaluations stored in MLE tables. Computing a dot-product reduces polynomials to a single value, i.e., the commitment.
MSMs are a bottleneck in ZKP provers and recent work has focused on accelerating them on ASIC and GPU (Daftardar et al., 2024; Zhang et al., 2021; Ma et al., 2023; Liu et al., 2024a; Ji et al., 2024; Lu et al., 2022; Zhou et al., 2024; Liu et al., 2024c, b; Hirner et al., 2025). The bottleneck is due to the extremely expensive elliptic curve point multiplications that they use. To reduce this cost, MSMs use Pippenger’s algorithm (Pippenger, 1976), which performs point multiplications via several point additions (PADDs). PADDs are still expensive, typically tens of regular modular multiplications. In the context of ASIC accelerators, the state-of-the-art (SZKP (Daftardar et al., 2024)) presents a framework for building scalable MSM architectures.
3. The HyperPlonk Protocol
In this section, we describe how HyperPlonk encodes the computation being proven, how this encoding shapes the structure of the underlying polynomials, and the key steps in proof generation.
3.1. Plonk-based encodings
In ZKPs, the program being proven must be converted into a specific form before generating the proof. While most prior works use Rank-1 Constraint System (R1CS) (lib, 2018; Groth, 2016)—a series of sparse matrix-vector encodings—HyperPlonk adopts a Plonk-based polynomial structure (Gabizon et al., 2019). R1CS and Plonk encodings map all program computations into addition, multiplication, and boolean operations, with nonlinear functions (e.g., branches) resolved via bit-wise decompositions.
Plonk-based encodings map each operation in a program’s execution to a gate – similar in principle to gates seen in other privacy-preserving protocols (Mo et al., 2023, 2025b; Chillotti et al., 2020; Feng et al., 2024) – that supports addition, multiplication, conditionality, and equality checks, as shown in Equation 1:
This can be viewed as a ZKP-specific instruction set where each gate acts as an instruction performing basic arithmetic or logical operations. In Equation 1, terms are binary control signals for left, right, multiply, and output ports. The term represents a constant input. Terms represent the gate’s inputs, and is its output. These last three terms are collectively called witnesses. An addition operation, for instance, is implemented by setting , , and other control and constant inputs to . This yields: . That is, if and only if the addition is correctly performed.
Each operation in a program is mapped to a Plonk gate. A program with operations has Plonk gates connected to form a circuit. A -variate polynomial represents the entire circuit, with binary assignments to variables representing individual gates. For example, if , the circuit has gates; represents the first gate.
is constructed using multilinear polynomials representing every term from Eq. 1. For example, is the left control input polynomial, is the first data input polynomial, and so on. Each of these polynomials also takes as an index. For example, in a program that maps to 16 gates, will tell us whether or not the multiplication control signal is enabled in gate 2. tells us what the output data signal is at gate 15. All individual polynomials are encoded in this fashion, and can be stored as MLE tables. These MLE tables are populated from the program trace in software before running the HyperPlonk prover. The MLEs are then used by the steps of the HyperPlonk protocol, which we discuss in the following sections.
3.2. SumCheck on Plonk-based Encodings
The key difference between the SumCheck used in HyperPlonk and the example in Section 2.2 lies in the structure of the underlying polynomials. While the earlier example operates on a single polynomial of arbitrary degree, HyperPlonk applies SumCheck to a version of the Plonk polynomial in Equation 1, as well as other polynomials discussed in Section 4.1. These polynomials consist of the products of multilinear polynomials, with terms of varying degrees. Figure 1 shows an example of how a SumCheck of the product of three multilinear polynomials is computed.
In the example from Section 2.2, if was multilinear, we would only need the evaluations at and because the result of summing over all variables yielded a univariate degree-1 polynomial. In this example, however, after summing over all variables, we are left with a univariate degree-3 polynomial. As mentioned in Section 2.3 (and from basic polynomial principles), a degree-3 polynomial must be evaluated at 4 unique points to fully characterize the polynomial. In the figure example, in the blue region, we are iterating over to compute the evaluations. For each iteration (red region), we must therefore evaluate each polynomial at . (Here, are the extensions that we compute from the existing evaluations at .) Then, in the green region, we compute the product for across all polynomials, and then sum across iterations. This is repeated for . After the summations, each polynomial’s MLE table is updated with a random challenge from the verifier. We refer to this step as MLE Update. For every boolean hypercube instance (e.g., in Figure 1), we extrapolate the linear polynomial in to the value of . For example, a table (for round 2) can be constructed from a table (from round 1) using the formula
The sum of all of these core computational steps is more expensive than baseline SumCheck. An additional layer of complexity is that HyperPlonk’s SumCheck polynomials exhibit degree variation across terms. This leads to imbalance in how many total evaluations are needed per term (i.e., needs 3 evaluations, while needs 4). This is handled by the HyperPlonk SumCheck protocol with a fixed interpolation step before the MLE tables are updated. While these computations are expensive, many of these operations can be performed in parallel. All polynomials can compute their extensions independently, and within a polynomial, extensions at different values can be parallelized. Thus, HyperPlonk’s SumChecks have high compute and degrees of parallelism.
SumCheck example illustrating the computation for an example polynomial.
HyperPlonk’s protocol steps, zKSpeed’s architecture, and dataflow for each protocol step.
Scenarios that can be done with Multifunction Tree.
3.3. HyperPlonk Protocol Steps
In this section, we outline the steps of the HyperPlonk protocol, as seen in Figure 2, what each step achieves, and what hardware units are needed to perform each step.
3.3.1. Witness Commits
The first step in HyperPlonk’s prover involves witness commitments. In this step, we compute MSMs between elliptic curve points and the witness polynomials to reduce each polynomial to individual commitments. The witness polynomials in HyperPlonk are typically “Sparse”, meaning that 90% of the values are either 0 or 1, and 10% are up to the full scalar bitwidth. SZKP (Daftardar et al., 2024) handles this by using Sparse MSMs. We implement Sparse MSMs using the MSM unit for this step.
3.3.2. Gate Identity
The Gate Identity step confirms that each gate in the circuit performs its computation correctly. The intuition is that if so, then in Equation 1 should evaluate to 0 for each gate, or equivalently, for each binary assignment of . Further, the sum of over all ’s should also equal 0, which can be confirmed by running SumCheck on . However, since the prior statement is only a necessary but not sufficient condition for correct computation, HyperPlonk performs SumCheck over a polynomial , where is a multilinear polynomial with evaluations computed from random challenges (the polynomial of Lemma 2.1 in (Chen et al., 2022)). We will henceforth refer to the construction of such a polynomial as Build MLE. Then, the sum over the boolean hypercube of should also be 0; this is known as a ZeroCheck. The Gate Identity step relies on three units: the Multifunction Tree unit (Section 4.3) to build , the SumCheck unit, and the MLE Update unit to update MLE tables between SumCheck rounds.
3.3.3. Wiring Identity
This step verifies that the outputs of each gate are routed correctly to inputs of downstream gates. This is achieved via a PermutationCheck (PermCheck), which is performed by constructing a numerator polynomial that encodes the natural ordering of the outputs and a denominator polynomial that encodes the permuted ordering (derived from permutation polynomials ), and then checking that , which we refer to as Fraction MLE, equals 1. As we will discuss, computing is time consuming, since inverting (or, in effect, performing divisions on) elements of a finite field is difficult. The Product MLE, , is generated by computing a cumulative product over to form the first half, then recursively applying cumulative products on itself to complete the second half. Then, the and polynomials are both committed to via MSM computations. A final ZeroCheck is also performed to ensure that the structure imposed upon and was not violated during their creation, i.e., the sum over the boolean hypercube of the resulting constraint polynomial should be zero everywhere. All polynomials are stored in the form of MLE tables. The Wiring Identity step uses the Construct N&D unit, the FracMLE unit, the Multifunction Tree unit (to construct the Product MLE and later for the Build MLE step within ZeroCheck), the MSM unit, the SumCheck unit, and the MLE Update unit.
3.3.4. Batch Evaluations
The SumCheck protocols require evaluating several MLEs at specific points. This is achieved with batch evaluations, where the input MLEs, as well as and , are queried at 6 points, some of which are derived from round challenges in the SumCheck portion of the Gate and Wiring Identity steps, and some of which are fixed at compile time. In total, 22 total evaluations are performed among 13 polynomials using 6 distinct points. This step is almost the exact reverse of the step in 3.3.2 used to construct , since we are compressing an entire MLE into 1 value, while is an entire MLE built from values. We will henceforth refer to the operation of evaluating an MLE at a point (consisting of field elements) as MLE Evaluate. These polynomial queries are then used by the verifier to check that the prover obeyed the protocol. The Batch Evaluation step uses the Multifunction Tree unit.
3.3.5. Polynomial Opening
This step succinctly verifies the correctness of the prover’s batch evaluations. For brevity, we will omit details of the polynomial opening step. To summarize, it involves first computing 6 MLEs as a linear combination of the MLEs from Eq. 1 and and . Then, 6 more MLEs are constructed from the query points in Section 3.3.4. These 12 MLEs are linearly combined into a single MLE, upon which a final SumCheck is computed. To avoid confusion, we will refer to this final SumCheck as OpenCheck and use SumCheck to refer to the underlying computation.
After OpenCheck, the first 6 MLEs used as OpenCheck’s inputs are linearly combined with the OpenCheck’s round challenges to construct a final MLE, which is denoted as . This MLE is first reduced to half its size, and it is used as the scalar set for a -point MSM. We then halve the scalar set and perform MSMs. For example, if , we compute a -point MSM, then a -point MSM, all the way to a -point MSM. The Polynomial Opening step uses the MLE Combine unit to compute the linear MLE combinations, the Multifunction Tree unit to build the 6 MLEs, the SumCheck unit to perform OpenCheck, and the MSM Unit to compute MSMs.
3.3.6. SHA3
zkSNARKs are non-interactive and use SHA3 to generate challenges as well as maintain transcripts. Transcripts are logs of proof-related computations checked by the verifier. In HyperPlonk, SHA3 is invoked between steps to “log” computed values, which is achieved by updating the SHA3 state. Intuitively, this ensures that all future challenges are bound to the existing history of computations logged in the transcript. As a result, SHA3 effectively acts as an order-enforcing mechanism. This means the protocol steps must be executed in series, as shown in Figure 2.
3.3.7. Compute Demands of HyperPlonk
Table 1 summarizes profiling results to characterize key functions and understand the sources of performance overhead and hardware needs when accelerating HyperPlonk. There are too many functions to list, so we present the twelve with the highest arithmetic intensity, which is defined as modular multiplications (modmuls) per byte, as done in prior work (Kim et al., 2022b; de Castro et al., 2021) (note SHA3 has no modmuls). Additionally, the reference CPU implementation is provided as a link. First, we observe that all functions require an immense amount of computation: ranging from millions to billions of 255/381-bit modmuls (comprising three integer multiplications) over all invocations of each function. For example, the data for Wire Identity MSMs modmuls reflects two function calls; for ZeroCheck Rounds there is one function call (see links in table). This motivates the need for both specialized modmul units and a high degree of parallelism to mitigate overhead. Second, compute intensity drops off sharply after the third function (since data reuse is limited), and the data input/output sizes for all functions are large, typically hundreds of megabytes up to terabytes as problem sizes scale to more gates. This motivates the need for large on-chip scratchpads to mitigate off-chip data movement when possible and high off-chip (i.e., HBM) bandwidth.
4. HyperPlonk Accelerator Units
HyperPlonk’s protocol is based on the BLS12-381 elliptical curve. Here, all MLE datatypes are 255 bits wide and all elliptical curve points in the MSMs are 381 bits wide. All MLE and MSM operations involve modular arithmetic primitives. These are built into each accelerator unit that requires them. zkSpeed comprises eight accelerator units and we describe each below.
4.1. SumCheck and MLE Update
Three HyperPlonk steps use SumCheck: ZeroCheck (Section 3.3.2), PermCheck (Section 3.3.3), and OpenCheck (Section 3.3.5). Each step has a unique polynomial shown, respectively, below:
In these equations, is a challenge from the SHA3 unit, and all other symbols represent multilinear polynomials. and are all Built MLEs constructed from SHA3 challenges (see the discussion of in Section 3.3.2). These SumCheck polynomials share a common sum-of-products representation, but are each unique and require slightly different datapaths. We develop a unified SumCheck unit that can handle each of these polynomials and highlight the key contributions of our architecture next.
4.1.1. SumCheck Round PE Microarchitecture
In ZeroCheck and PermCheck, there are polynomials that appear multiple times across terms. In HyperPlonk’s CPU baseline, the boolean hypercube summations are performed iteratively term-by-term, incurring redundant computation for these repeating polynomials. We address this by computing all evaluations for each polynomial in parallel. For example, in Equation 3, the polynomial has to be evaluated at and because the largest-degree term it appears in has a degree of four. This computation only needs to be performed once before being used to compute the product in each term of . Figure 4 shows an example of how our SumCheck PEs handle ZeroCheck computation for . Each unique polynomial takes its respective values (e.g., in Figure 4), and uses these values to extend to the needed evaluations. Then, each product in Equation 3 must be computed for each evaluation. The products at the evaluation points are then accumulated into registers. Due to the inherent degree imbalance, some terms have fewer evaluations; the additional evaluations are computed via Barycentric Interpolation (Berrut and Trefethen, 2004). This is omitted in the figure since it only adds a fixed cost at the end of each round (23 modmuls for ZeroCheck and 46 modmuls for PermCheck). Each polynomial observes a different datapath, so we use a specialized design to exploit high reuse, full-pipelining, and high levels of parallelism.
4.1.2. Streaming approach
At the start of the protocol, the MLE tables corresponding to polynomials in Equation 1 are all provided to the prover and can be stored on-chip. However, as these MLEs undergo rounds of SumCheck, the process of incorporating challenges into the MLE tables expands binary values to full 255-bit values. Though the number of MLE table entries reduces by half each round, the data itself grows by over 100 between rounds 1 and 2, so the total storage cost for storing all MLEs is intractable. However, each round, the intermediate values of MLE tables are only used by the main SumCheck computation and then by the MLE Update unit to be halved in size. Since there is no data reuse in-between rounds, we adopt a streaming-based solution to alleviate the pressure on on-chip SRAM storage. The key tradeoff here is that our SumCheck and MLE Update units become memory-bound, since each MLE in Equations 3-5 must be updated and written back to off-chip memory after SumCheck rounds, necessitating off-chip traffic. Fortunately, recent advances in high-bandwidth memory (HBM) can supply very high bandwidths to offset this. This is inline with many other cryptographic computing accelerators, which also rely on HBM (Daftardar et al., 2024; Ebel and Reagen, 2024; Samardzic et al., 2021, 2022; Kim et al., 2022b, 2023, a). We analyze the bandwidth sensitivity of the SumCheck computations in our evaluation section.
4.1.3. Scaling to Multiple PEs
Each SumCheck PE handles the product and sum for one iteration over the boolean hypercube. For example, in Figure 1, this corresponds to one red region per polynomial (MLE table) and the first product in the green region for each . These iterations run in parallel with multiple PEs storing their own accumulation registers. The bottom of Figure 4 shows the accumulation of each evaluation across boolean hypercube values (equivalently, across indices of the MLE tables). After all evaluations are done, an MLE Update PE handles the updates for one MLE table and can provision multiple modmuls. Multiple PEs can run in parallel, handling MLE tables independently.
4.1.4. Unified SumCheck PE
The three polynomials , , have different datapaths. We use HLS tools to generate a unified PE that handles each SumCheck variation used in HyperPlonk. Each PE requires 94 modular multipliers, compared to 184 modular multipliers without resource sharing, saving 48.9% on area.
4.1.5. Comparison to Prior SumCheck Implementations
HyperPlonk’s CPU library is designed to support any composition of multilinear polynomials for different protocols, not just the three we have shown. Consequently, repeating polynomial computations as observed in the CPU implementation greatly improves programmability, as opposed to having specialized functions to handle the specific computation patterns we optimize for.
NoCap (Samardzic et al., 2024) is a recent accelerator that also accelerates SumCheck, but there are critical differences at the protocol-level that motivate our architecture. NoCap implements Spartan (Setty, 2020), which uses R1CS encodings resulting in two SumCheck instances that look as follows: and . NoCap uses a vector architecture with 2048 PEs to process boolean hypercube instances with a Beneš network to sum across PEs. This makes sense for NoCap because Spartan’s polynomials are degree and with up to two terms. In contrast, HyperPlonk’s polynomials have a more heterogeneous structure; there are more terms of varying degree. This complexity arises from the usage of the control and constant polynomials () to represent gates. These are required to keep HyperPlonk verifier costs low (in Section 7, we see NoCap’s verifier is slower). Consequently, mapping HyperPlonk’s polynomials to a vector architecture would put more pressure on vector register files because of the high amount of intermediate values needed to be read. As seen in Figure 4, our SumChecks also require very complex communication, which can increase bandwidth pressure and may not be efficient to implement with a Beneš network. Further, our specialized PEs immediately reuse values without relying on register files to store the numerous amount of intermediates. We have fewer, heftier PEs, so data movement costs are less relative to those incurred by a vector processor.
4.2. MSMs
PADD bucket.
MSMs are used in three HyperPlonk steps: Witness Commits, Wiring Identity, and Polynomial Opening. Several prior works accelerated MSMs; we begin with the base MSM design from SZKP (Daftardar et al., 2024), the state-of-the-art for accelerating Groth16, and propose two optimizations for better performance and area efficiency.
In HyperPlonk, the Sparse MSMs run in series and are on the critical path. This is unlike prior protocols like Groth16 where their execution could be masked via parallel processing with the Dense MSMs. Consequently, we opt to use the same MSM hardware unit for both Sparse and Dense MSMs in zkSpeed (as opposed to separate units in SZKP). zkSpeed uses a similar scheme as SZKP for handling sparse computations. First, we compute the sum of all points corresponding to -valued scalars. This is done by fetching the points corresponding to -valued scalars into the MSM unit’s SRAM banks. Then, using a tree-based approach, we feed two points at a time to the pipelined Point Adder (PADD) unit, with the result written back to the SRAM banks. This process repeats until we have reduced all points to the final sum for points corresponding to -valued scalars. Note that in this step, we need not fetch scalars into the MSM’s SRAM banks since they are all . Then, we use Pippenger’s algorithm (Pippenger, 1976) on the remaining 10% of (dense) scalars. For the Dense MSMs in the Wiring Identity and Polynomial Opening steps, we use Pippenger’s algorithm for the full MSM computation. We now discuss the two improvements to SZKP’s MSM architecture.
4.2.1. Reduced Memory Footprint
First, we note that elliptical curve points, while being three-dimensional, are initialized as , 1) coordinates in HyperPlonk. Thus, we only fetch two coordinates per point, saving off-chip bandwidth. We further save on-chip SRAM area. SZKP provisions one scalar memory bank and three point memory banks, to hold coordinates. While Sparse MSMs need not store the 1-valued scalars, the tree-based addition partials still need buffering (their coordinates are no longer 1-valued) necessitating the coordinate memory bank. Therefore, zkSpeed allocates three SRAM memory banks. In Dense MSM operation, the memory bank is reused to store the scalars, and since partial sums are only stored in bucket registers, we do not need to provision another dedicated memory bank, as in SZKP. This represents a savings of 18% in on-chip SRAM area compared to having a dedicated scalar memory bank.
4.2.2. Faster Bucket Aggregation
The second optimization addresses a runtime bottleneck in SZKP’s bucket aggregation step. After sorting points into buckets and computing each bucket’s sum, SZKP employs a naive aggregation algorithm to calculate the sum , where is the window size (Daftardar et al., 2024; Pippenger, 1976), and represents the accumulated sum of the -th bucket. This is inefficient when processing smaller MSMs, e.g., 32-point MSMs, which are prominent in Polynomial Opening. The fixed bucket aggregation latency becomes a performance bottleneck because the point additions are serially performed and do not leverage the pipelining available in the PADD unit. Consequently, the PADD is underutilized in this step. To address this, zkSpeed adapts bucket aggregation introduced in (Liu et al., 2024a). This scheme divides aggregation into smaller groups, computes the partial sums within each group in parallel, and finally combines the results. As shown in Figure 5, it reduces the bucket aggregation latency by an average of 92% across all window sizes compared to SZKP. We select a group size of 16, which provides the best overall performance and ensures the aggregation step no longer dominates runtime for small MSMs.
4.3. Multifunction Tree Unit
4.3.1. Tree-like Computation in HyperPlonk
Many HyperPlonk functions exhibit binary-tree compute patterns, including Build MLE, MLE Evaluate, and constructing Product MLE (). zkSpeed supports these in hardware with our Multifunction Tree Unit (MTU), specially designed to handle these compute patterns effectively. Build MLE is a function used in ZeroCheck and OpenCheck steps. It constructs a table with entries from random values to , where represents problem size. The computation is divided into layers (as a binary tree) to reduce the number of modular multiplications from to . The multiplier tree is used for batch inversion during Fraction MLE generation (see Section 4.4), where it efficiently calculates the product for an inversion batch size, . Similarly, the MLE Evaluate in the Batch Evaluation step operates like a multiplier tree but includes additional modular additions in each operation. The Product MLE generation in the Wiring Identity outputs all layer results. The functions’ dataflows are presented in Figure 3.
4.3.2. zkSpeed’s Tree Approach
The original (CPU) HyperPlonk implementation uses breadth-first (level-order) traversal (BFS). This is inappropriate for hardware acceleration as it puts increased pressure on SRAM capacity and off-chip bandwidth. For example, a problem size of requires up to intermediate elements in a level, each 255 bits wide, which would require 128MB for the intermediates alone. We propose a hybrid strategy that applies depth-first traversal (DFS) to the upper levels of the binary tree and switches to BFS at lower levels. This approach leverages efficient memory reuse in the upper levels through DFS, while enabling parallelism in the lower levels via BFS. It reuses and consumes intermediate results as they are produced, reducing the amount of data that must be stored on-chip or spilled to DRAM.
Hardware structure of the Multifunction Tree unit.
4.3.3. Multifunction Tree Architecture
We further analyze this tree architecture in (Mo et al., 2025a). Figure 6 shows an example of how the MTU works. Each PE includes a modular multiplier and modular adder444For Build MLE, and only require one modular multiplication since is computed as . The hardware supports three processing modes of Figure 3. For the Inverse Tree (e.g., MLE Evaluate), the unit accepts inputs in parallel and reduces tree-level partials via a matching hardware dataflow; in Figure 6, , and the data flows from left to right. If the tree has more levels than hardware supports (three levels in the example; a problem size workload has 20 levels), the remaining levels are scheduled and processed in DFS via the accumulator. Outputs from the hardware tree (Level 3 in Figure 6) are pushed in an accumulator-local register file and once operands are ready, popped to the accumulator PE. The bottom of Figure 6 shows its schedule. Initially, there are gaps as the accumulator must wait two cycles for each input pair. However, once multiple levels are processed, the gaps are filled; this can be seen in cycle 44, where levels 4 and 5 are processed by the PE at the same time (i.e., both in the pipeline). Thus the PEs in the tree have high utilization: for a workload, they are over 99% utilized during the whole computation. The red arrows coming out of each PE show how computations are sampled and output to create the Product MLE. The Forward Tree (Build MLE) support is shown in blue, and data flows from right to left. Each PE takes the previous level and a challenge as inputs and generates two outputs that are fed to the next level. Similarly, if the tree has more levels, the accumulator PE is scheduled to generate (roughly) one output per cycle to feed the rest of the PEs, corresponding to the last 3 levels, outputting 8 results in the end. Switching is supported by muxes at PE inputs.
The advantages of this hybrid traversal (Mo et al., 2025a) are noticeable: it eliminates the need to store entire intermediate levels, making it practical for large problem sizes, and provides the ability to rate-match with upstream or downstream units and maintain throughput. By adjusting the number of PEs, the unit can handle varying input and output rates, forming a full pipeline with other units. The ability to reuse across multi-functions eliminates the need to allocate multiple dedicated units, saving 41.6% area across global Pareto design points in Section 7. As HyperPlonk’s code uses BFS, it experiences greater dependence distances. Our traversal mitigates this by scheduling work to avoid dependence stalls. Executing nodes already stresses CPU’s limited compute resources, so we expect hybrid traversal to have little performance impact in software.
4.4. Construct N&D and FracMLE
4.4.1. Creating with Constant-Time Inversion
The Construct N&D stage generates the and MLEs discussed in Section 3.3.3. Elements of six intermediate MLEs, , and , are computed in parallel from modular additions and multiplications of the witness () and wiring permutation () MLEs stored in on-chip SRAM, and two challenges from SHA3. These intermediate MLEs are written off-chip for the subsequent PermCheck and multiplied to obtain the and (e.g., ) elements, and fed to the FracMLE unit.
The Fraction MLE, , requires computing the modular inverse of every element of the Denominator MLE (), and multiplying each inverted element with the corresponding element of the Numerator MLE (). Given , modular inversion outputs a , such that . We use the constant time Binary Extended Euclidean Algorithm (BEEA) (Pornin, 2020) for this operation to make it data-oblivious and ensure constant scheduling. The algorithm requires iterations of a loop consisting of shifts and subtractions, where is the number of input bits (), resulting in a 509-cycle latency. A data-dependent implementation is faster for smaller input values, but since we are computing on random inputs (derived from SHA3 hashes), the average latency of such an implementation (2 cycles) is only 1 better. In exchange for this negligible latency overhead, our constant-time implementation reduces design complexity as this ensures elements of (and thus ) are generated in-order if we execute multiple inverses in parallel (as we will discuss shortly). Parallel execution of the data-dependent algorithm may output elements out-of-order and would require buffering or stalling to resolve.
4.4.2. Batching for Modular Inversion
Computing modular inverses for each 255-bit MLE element is expensive, so we heavily optimize it. We leverage Montgomery batching (Montgomery, 1987) to amortize the cost of one inversion across multiple elements and improve per-inversion throughput. In the standard approach (also implemented in (Chen et al., 2022)), to compute the inverses of a batch of elements (e.g., ), partial products are computed sequentially (e.g., , , , ) and the final product is inverted. This single inverse is then propagated backward to recover the individual inverses (, etc.) through additional sequential multiplications. While this reduces the number of inversions, the sequential multiplications and high modular multiplication latency can limit hardware performance.
We address these limitations with two modifications. First, we use a multiplier tree, detailed in Section 4.3; this significantly reduces latency and improves scaling for large values of the batch size, which we denote , from to multiplications. Second, we overlap the sequential multiplications for partial products with modular inversion to mask long multiplication latencies.
4.4.3. Batch Inversion Architecture
Figure 7 illustrates the architecture of our modular inverse unit using batching. We use multiple batched inversion units in round-robin fashion to completely mask long inversion latencies and enable the FracMLE unit to accept one input and generate one output per cycle, behaving as a pipeline with depth where is the number of FracMLE units. We achieve this by using enough batched inverse units to mask the latency of one batch inversion. The multiplier tree and one multiplier are reused across all units to compute their inputs and produce the individual elements of , respectively. One batch inversion latency is the maximum of the parallel partial product latency and the combined multiplier tree and modular inverse latencies. The former scales while the latter scales .
4.4.4. Optimizing for Batch Size
We frame the choice of batch size as an optimization problem to minimize latency imbalance (between partial products and the multiplier tree and inversion) and area. As shown in Figure 8, the left y-axis plots latency imbalance, which is initially high due to the fixed cost of modular inversion. It decreases as partial product latency rises, reaching a minimum at , then increases again as partial-product latency begins to dominate. The right y-axis shows total area (including multiplier trees and partial products SRAM), which also reaches a minimum at due to fewer required inverse units (e.g., 256 units for vs. only 12 for ). Additionally, starting at , we can reuse the multiplier tree across all units since its latency allows it to complete one batch before the next arrives. Beyond , partial-product latency surpasses modular inversion, and SRAM overheads continue to grow without reducing the number of inverse units—diminishing returns in both performance and area. Based on these tradeoffs, we select as the optimal batch size.
4.5. MLE Combine Unit
The MLE Combine Unit is used in the Polynomial Opening step. As mentioned in Section 3.3.5, there are several linear combinations of MLEs that are performed before OpenCheck and before the MSMs. These operations are straightforward, and use a combination of MLEs stored in on-chip SRAM and off-chip memory to construct the MLEs used in subsequent steps. Because OpenCheck happens in series with the MSMs, the respective MLE Combine operations also happen in sequence. Consequently, we can share resources between these two operations. For the design point we highlight in Table 5, without sharing, we would require 122 modular multipliers. With sharing, we require only 72, representing 41% area savings.
4.6. On-Chip MLE SRAM
The MLE table size scales with the number of gates, i.e., for a problem size of gates, each MLE table has entries. In practice, input MLEs are sparse. Control MLEs are all binary, and are roughly 90% 1s and 0s and 10% full bit-width. These MLEs get reused throughout the protocol, so we store them on-chip in global SRAM. We compress the tables, packing together control MLEs and using address translation units to perform lookups to either binary or 255-bit data. These compression strategies save to on MLE storage across problem sizes. This also significantly reduces off-chip accesses; for example, in the Polynomial Opening step, only 2 of 13 MLE tables ( and ) are stored off-chip, cutting bandwidth usage by 84%.
5. zkSpeed Architecture HyperPlonk Mapping
Figure 2 provides an overview of the HyperPlonk protocol and zkSpeed architecture. Colors indicate the mapping of protocol steps (Section 3.3) to the accelerator units (Section 4) they run on. The zkSpeed architecture is streaming in nature and captures on-chip data reuse when feasible via explicitly managed scratchpad memories. The architecture has four major components: accelerator units, local and global SRAM, a multi-channel shared bus, and HBM interface. HBM is needed to feed the chip with high bandwidths needed by HyperPlonk, and we conduct bandwidth sensitivity studies in Section 7.
Dataflow. HyperPlonk is data oblivious at the stage granularity. This allows zkSpeed to statically schedule computations and manage units, SRAM, and the buses via a simple controller. Further, zkSpeed uses a shared bus rather than a crossbar or NoC. This design choice was made after rigorously analyzing the HyperPlonk dataflow (see Figure 2C). We observed that at any given time, only 1-2 zkSpeed units typically communicate, and at most 4 independent bus channels are needed to avoid stalling – this is during Wire Identity where Construct N&D sends results to FracMLE, FracMLE simultaneously feeds ProdMLE and MSM units, and ProdMLE streams to MSM. Units overlap computation with each other, e.g., enabling MSM to start processing partial outputs from FracMLE, effectively masking latency. Without bus stalls, we are able to rate match each accelerator unit to pipeline across modules when possible, further improving performance.
zkSpeed deploys a highly banked global SRAM and two local SRAMs for FracMLE and MSM units that store data unused by other units (the FracMLE SRAM captures MLE table reuse, the MSM’s SRAM reuses elliptical curve points). All other units share the global SRAM, which stores input MLEs. At the start of execution, these MLEs are prefetched from HBM and remain unchanged on-chip throughout execution. They are read at the beginning of multiple protocol steps, thereby reducing HBM pressure. zkSpeed allocates a single-channel shared bus for units to read the global SRAM since only one unit requires access at any given time. HBM access is managed by a memory controller that has dedicated point-to-point connections to each module and the global SRAM, with enough wires to a given component to accommodate the widest access needed. The controller interfaces with the HBM PHYs and arbitrates access to the HBM channels to ensure that no channel is being used more than once simultaneously (i.e., no channel conflict).
zkSpeed Programmability. zkSpeed modules are programmed by instructions specifying problem sizes and configuring local controllers and address data from the buses, SRAM, and HBM. Due to the ASIC nature, much of the fine-grained control is handled by FSMs within each unit. For each HyperPlonk protocol step, each unit (e.g., bus, modules, global SRAM, and HBM controller) are configured with complex instructions and run to completion. Then next set of instructions are loaded to execute the next protocol step.
HyperPlonk Trends and Outlook. Zero-Knowledge protocols are still continuously evolving and improving in several aspects. However, zkSpeed, with its focus on HyperPlonk, still has significant stability. Firstly, HyperPlonk has seen adoption both in industry implementations (Zhang et al., 2022; Diamond and Posen, 2023) and academic research (Kohrita and Towa, 2024; Diamond and Posen, 2023, 2024). In addition, our design is modular, and the key components of HyperPlonk SumChecks, MSMs, and MLEs, are present in essentially any modern SNARK protocol (Thaler, 2013; Setty, 2020; Groth, 2016; Setty et al., 2024; Kothapalli et al., 2022; Bünz and Chen, 2023). Therefore, zkSpeed can also be targeted to new protocols so long as they comprise the fabricated modules, including proof composition methods (Valiant, 2008) that seek to compose protocols like Orion and HyperPlonk.
6. Methodology
6.1. Performance Modeling
SumCheck has a fixed, data-oblivious dataflow, which allows us to model its performance analytically. Modules that feed into the SumCheck units are also data-oblivious and are modeled similarly, accounting for rate-matching assumptions. For the MSM, we use a cycle-accurate simulator to model performance. Each unit within zkSpeed is modularized. To understand full-chip performance, we conduct a design space exploration of all combinations of design parameters detailed in Table 2 and then analyze the pareto-optimal space to pick a suitable configuration for profiling runtimes. We also construct power traces to estimate average power for the full-chip architecture. We use Catapult HLS 2023 to generate the RTL for Montgomery multipliers (as done in prior work (Daftardar et al., 2024)), the fully-pipelined, unified SumCheck unit (to handle ZeroCheck, PermCheck, and OpenCheck), and the fully-pipelined PADD unit. Consistent with HyperPlonk, we use the BLS12-381 elliptical curve, where all MLE datatypes (e.g., in the SumCheck unit) are 255-bit, and all elliptic curve points (e.g., in the PADD) are 381-bit. Using Design Compiler with TSMC 22nm, we find the critical path in our design is the 381-bit PADD unit at 1.05ns. We use Synopsys 22nm Memory Compiler to generate SRAM estimates. For SHA3, we use the IP block from OpenCores (OpenCores, 2013). We scale down to 7nm using scale factors of 3.6 for area, 3.3 for power, and 1.7 for delay (as in prior work (Daftardar et al., 2024)), and clock all zkSpeed accelerators at 1 GHz.
6.2. Benchmarks
HyperPlonk was evaluated using mock circuit workloads (Chen et al., 2022), as there is no publicly available compiler to generate real workloads. We similarly use synthetically generated workloads to model the performance of our architecture, which is standard for ZKP benchmarks, as performance primarily depends on the size of the workload. Similarly, GZKP (Ma et al., 2023) uses synthetic workloads to benchmark their implementation for workload sizes of and higher. NoCap (Samardzic et al., 2024) uses workloads from libsnark (lib, 2018) that prove relatively small circuits and scales them up to larger problem sizes because their proving times are dominated by fixed overheads on smaller ZKP circuits. In the context of HyperPlonk, workload statistics primarily affect the witness commit (Sparse MSM) step. The scalar distributions of the MSMs in the Wiring Identity and Polynomial Opening are random because they are constructed in part from SHA3 challenges. Therefore, while all MSMs are data-dependent, the overall runtimes are effectively workload-agnostic and roughly the same at iso-problem size. All other steps in HyperPlonk are data-oblivious. From prior work (Daftardar et al., 2024; Ma et al., 2023; Zhang et al., 2021; lib, 2018), we know that in Sparse MSMs, the scalars are typically 5-10% dense (i.e., full bit-width), and 95-99% sparse. Since the dense components of Sparse MSMs are still runtime dominant, we assume a pessimistic upper bound of 10% dense scalars and 90% sparse scalars, of which 45% are 1s and 45% are 0s. The rest of the protocol steps operate on 100% dense scalars/MLE values. In our evaluation, we use five workloads from prior work (Chen et al., 2022; lib, 2018), and show our results in Table 3.
7. Evaluation
7.1. Pareto Space Analysis
Figure 9 shows the design space for a problem size of gates under seven bandwidth scenarios. We sweep all the parameters in Table 2 and obtain Pareto curves for each bandwidth individually, and then construct the global Pareto curve from these seven local Pareto curves. The key highlight from this plot is that HBM3-scale bandwidths (e.g., 1-4 TB/s (hbm, 2023)) do yield significant performance gains over, e.g., an HBM2-scale bandwidth (0.5 TB/s (Inc., 2020)). Beyond 300 mm2, the globally Pareto-optimal design configurations yield over 2 speedups compared to 512 GB/s designs and over 700 speedups over the CPU baseline. This is because high-performance SumCheck designs quickly saturate 512 GB/s of bandwidth. In this analysis, we also include the cost of HBM PHYs (Samardzic et al., 2024; Ebel and Reagen, 2024; Mo et al., 2023; Inc., 2020; Samardzic et al., 2021, 2022; Kim et al., 2022a, b, 2023) where the PHY cost is 14.9 mm2 for a single HBM2 PHY and 29.6 mm2 for a single HBM3 PHY. For low-performance designs (below 100mm2), high-bandwidth memory becomes less effective, as these designs tend to underutilize the available bandwidth while still incurring the high PHY area overhead associated with HBM-scale memory. Figure 9 shows that zkSpeed remains viable even at bandwidths typical of DDR5 (Schlachter and Drake, 2019) (256 GB/s and below). While HBM allows exploring scalability benefits, less expensive memory technologies can be used to achieve Pareto-optimality within a target performance range (e.g., within 50 ms).
Pareto Frontiers for Gates. We plot both the individual Pareto curves for each bandwidth model, and then the globally optimal Pareto curve across all bandwidths in the inset.
Area breakdown and runtime for the seven Pareto points we choose.
| Module | Design Knob | Values |
| MSM | Cores | 1, 2 |
| MSM | PEs per Core | 1, 2, 4, 8, 16 |
| MSM | Window Size | 7, 8, 9, 10 |
| MSM | Points/PE | 1K, 2K, 4K, 8K, 16K |
| FracMLE | PEs | 1, 2, 4 |
| SumCheck | PEs | 1, 2, 4, 8, 16 |
| MLE Update | PEs | 1, 2, , 11 |
| MLE Update | Modmuls/PE | 1, 2, 4, 8, 16 |
| Memory | Bandwidth (GB/s) | 64, 128, 256, 512, 1T, 2T, 4T |
We analyze the area and runtime of selected points on the Pareto curve in Figure 10 to further understand bandwidth sensitivity. We pick Pareto points representing the highest-performing design point for each bandwidth level. In the area breakdown, moving from low to high-performance design points (A to D), the proportion of the SumCheck area increases significantly, because SumCheck is bandwidth-intensive, and higher bandwidth allows more parallel SumCheck PEs, boosting throughput. The MSM unit accounts for a large portion of the total area, but its absolute area remains unchanged when switching to high performance. This trend is also evident in the runtime breakdown: total runtime decreases as bandwidth increases, and the runtime contributions of the SumCheck-related processes (ZeroCheck, PermCheck, and OpenCheck) become smaller. Our analysis of the Pareto design points shows that high performance significantly depends on sufficient bandwidth, particularly improving the SumCheck computation. Conversely, for low-performance designs, the system utilizes less bandwidth and allocates more resources to MSM computation.
7.2. Bandwidth Sensitivity
Figure 11 shows how the speedups for MSM-related computation and SumCheck-related computation scale with increased PE count and bandwidth. These are two runtime-dominant components in our design points shown in Figure 10. We take the runtime of all MSM and SumCheck operations for 1 PE under 512 GB/s, and compute respective speedups to these numbers. Because MSMs are compute-bound, adding compute resources improves the speedups significantly, while adding bandwidth does not. We do not see perfectly linear speedup because of the serialization incurred in Polynomial Opening MSMs. SumChecks, which rely on a streaming-based approach, are memory-bound. As we add compute, we see linear speedups initially and then diminishing returns after saturating bandwidth. Consequently, in our Pareto-optimal design space, we see most points along the curve outside high-performance regimes use at most 2 SumCheck PEs compared to 8 or 16 MSM PEs.
7.3. CPU Performance Comparisons
Our CPU is an AMD EPYC 7502 32-core processor (Morgan, 2019a; AMD, 2024; tech, 2024; Daftardar et al., 2024). The total die size is 296 mm2. We sweep the problem sizes, and for each problem size, pick a Pareto-optimal design point that is close to 296 mm2. In these comparisons, we exclude the PHY cost, since the AMD EPYC processor has its own separate die for I/O (Morgan, 2019b). Therefore, we compare our total compute and on-chip memory area with the CPU’s total core area, including on-chip caches. We assume 2 TB/s HBM to achieve Pareto-optimality in Figure 9.
7.3.1. Runtime Breakdown and Utilization
Figure 12 shows the latency breakdown of HyperPlonk proving on a CPU and a pareto-optimal zkSpeed design for a problem size of gates. The CPU executes kernels sequentially, enabling detailed profiling; we report step latency for zkSpeed due to its parallel scheduling of kernels. As expected, the majority of time goes to processing MSMs, while a handful of other kernels account for single percentage points of runtime. Figure 13 presents the utilization of each unit and relative (datapath) area allocation (design in Table 5). The utilizations vary from over 70% to 5% for some modules. zkSpeed was intentionally designed (via the design space search in Figure 9) to allocate resources to cores to optimize high performance per area, i.e. the Pareto front. This can be seen in our analysis in two ways. First, the cores taking up most area, notably MSM at 64.6%, are the most used, and following the profiling data (Figure 12) require the most speedup. Second, though some units are used infrequently, they (i) take up little area and (ii) are essential to accelerate to achieve the speedups desired (i.e., 2-3 orders of magnitude). For example, the SHA-3 unit is rarely used, but provides a speedup of over over the CPU and takes only area. Additionally, consider that MLE Combine makes up 3.3% of the CPU runtime, but without acceleration caps speedup to a mere 30.3, thus justifying its 5.85% area allocation and relatively low utilization.
run time breakdown.
utilization breakdown.
7.3.2. Iso-CPU Area Comparisons
After picking each Pareto-optimal design, we run synthetic benchmarks over problem sizes . Figure 14 shows the speedup of each design over CPU baseline, and the breakdown across different steps of the protocol to understand where our speedups come from. In general, we get more speedup from our MSM units than the SumCheck units. This intuitively follows, given our observations that MSMs are compute-bound and more robust to bandwidth constraints. Additionally, the CPU poorly handles sparse computations because it serially computes the point addition for -valued scalars. Polynomial Opening MSMs also incur serialization costs that we reduce by overlapping MSM executions where possible. The variations in speedups over different problem sizes are an artifact of our choice to highlight different Pareto points per problem size; for example, at size problems, a dual-core MSM is used, while at size a single-core MSM is chosen. This is because on-chip MLE SRAM area begins to dominate, limiting MSM compute area and therefore achievable speedup. Storing MLE tables entirely off-chip may improve MSM speedups at higher SumCheck bandwidth costs. These tradeoffs can be explored in future work.
| Workload |
Runtime (ms) | |||
| CPU | zkSpeed | ||||
| Zcash | 1429 | 1.984 (720) | |||
| Auction | 8619 | 11.405 (755) | |||
| Rescue-Hash Invocations | 18637 | 22.082 (844) | |||
| Zexe’s Recursive Circuit | 37469 | 43.451 (862) | |||
| Rollup of 10 Pvt Tx | 74052 | 86.181 (859) | |||
| Accelerator | NoCap | SZKP+ | zkSpeed |
| Protocol | Spartan+Orion | Groth16 | HyperPlonk |
| Main Kernels | N & S | N & M | S & M |
| Encoding | R1CS | R1CS | Plonk |
| Proof Size | 8.1 MB | 0.18 KB | 5.09 KB |
| Setup | none | circuit-specific | universal |
| Prime | fixed | arbitrary | arbitrary |
| Bit-width | 64 | 255b/381b | 255b/381b |
| CPU Prover (s) | 94.2 | 51.18 | 145.5 |
| HW Prover (ms) | 151.3 | 28.43 | 171.61 |
| Verifier (ms) | 134 | 4.2 | 26 |
| Chip Area (mm2) | 38.73 | 353.2 | 366.46 |
| # Modmuls | 2432 | 1720 | 1206 |
| Modmul (mm2) | 0.002 | 0.133 / 0.314 | 0.133 / 0.314 |
| Power (W) | 62 | ¿220 W | 170.88 |
7.4. Workload Evaluation
We pick a fixed design and show the end-to-end speedups in Table 3. As mentioned in Section 6.2, we assume pessimistic 10% sparse scalar statistics for each workload. Our design has one MSM unit with 9-bit windows, 16 PEs, and 2048 points per PE, with 1 FracMLE PE, 2 SumCheck PEs, 11 MLE Update PEs, and 4 modular multipliers per MLE Update PE. The area breakdown is provided in Table 5. At roughly iso-CPU-core compute cost, zkSpeed achieves a geometric mean speedup of 801 over the CPU, with total area of 366.46 mm2, total average power of 170.88 W, and total power density of 0.46 W/mm2, which is within that of our CPU (Daftardar et al., 2024).
| Module | Area (mm²) | Average Power (W) |
| MSM (16 PEs) | 105.64 | 76.19 |
| SumCheck (2 PEs) | 24.96 | 5.38 |
| Construct N&D | 1.35 | 0.19 |
| FracMLE | 1.92 | 0.25 |
| MLE Combine | 9.56 | 0.34 |
| MLE Update | 5.84 | 1.13 |
| Multifunction Tree | 12.28 | 4.16 |
| Other | 1.98 | 0.04 |
| Total Compute | 163.53 | 87.68 |
| SRAM | 143.73 | 19.60 |
| HBM3 (2 PHYs) | 59.20 | 63.60 |
| Total Memory | 202.93 | 83.20 |
| Total | 366.46 | 170.88 |
8. Related Work
Much of the prior body of cryptographic hardware and systems research has focused on Fully Homomorphic Encryption and Multi-Party Computation (Kim et al., 2022a, b, 2023; Samardzic et al., 2021, 2022; Soni et al., 2023; Mo et al., 2023; Garimella et al., 2023; Neda et al., 2024; Ebel et al., 2025; Jha and Reagen, 2024). ZKP hardware research is relatively newer, and has focused primarily on accelerating NTTs and MSMs (Liu et al., 2024a; Ji et al., 2024; Lu et al., 2022; Liu et al., 2024c; Zhou et al., 2024; Liu et al., 2024b; Qiu et al., 2024; Butt et al., 2024; Zhu et al., 2024; Chen et al., 2024; Wang and Gao, 2023; Yang et al., 2025; Wang and Gao, 2025; Hirner et al., 2025; Zhang and Franchetti, 2025). A few works have accelerated SumChecks on GPU (Lu et al., 2024) and ASIC (Samardzic et al., 2024) as well as hashing alternatives to SHA-based hash functions (Sheybani et al., 2025; Ahmed et al., 2024; Wang and Gao, 2025). Some systems accelerate end-to-end Groth16 proofs (using NTTs and MSMs) on GPU (Ma et al., 2023) and ASICs (Daftardar et al., 2024; Zhang et al., 2021), while others accelerate Spartan proofs with Orion commitments (using NTTs and SumChecks) (Samardzic et al., 2024). We compare zkSpeed with two ASICs, SZKP and NoCap, that accelerate full proofs end-to-end.
SZKP is the state-of-the-art for accelerating Groth16 proofs, focusing on scalable MSM designs and (quasi)-deterministic scheduling for Pippenger’s algorithm. It accelerates all MSMs, including Sparse G2 MSMs (not present in HyperPlonk), achieving geomean speedups of 493 over a CPU. SZKP improves on PipeZK (Zhang et al., 2021), the first hardware accelerator for Groth16 proofs. While Groth16 and HyperPlonk have similar application spaces, as mentioned in Section 1, the key advantage of using HyperPlonk is the universal setup, which means that the protocol parameters are application-agnostic. For Groth16, every new application that wants to use a ZKP needs its own trusted setup ceremony (cer, 2023), which is impractical as the application space grows. Given this context and the recent shift away from Groth16 (Nelson, 2022), the slightly larger proof sizes are considered a reasonable tradeoff.
NoCap is a vector-based processor for accelerating Spartan+Orion proofs, but its application space differs from zkSpeed’s. NoCap thrives in applications where proof size is not critical, or there are few verifiers. It achieves geomean speedups over PipeZK. In contrast, zkSpeed is ideal for many verifiers and in consensus-based systems; this is where ZKPs are experiencing growing interest.
Comparison: Table 4 compares zkSpeed, NoCap, and SZKP’s protocols, software, and hardware costs. zkSpeed’s parent HyperPlonk has the slowest software prover, reflecting the complexity of the protocol. Of note, Spartan’s prover is slow; NoCap’s authors explain this is due to inefficient implementation.
We compare NoCap’s hardware implementation using the design point and numbers from their paper scaled to 7nm using scale factors from prior work (Mo et al., 2023; Daftardar et al., 2024). We then select a zkSpeed configuration with roughly similar prover time to NoCap. At iso-prover time, zkSpeed incurs a nearly area cost in return for a three orders-of-magnitude reduction in proof size. NoCap’s lower costs come from not having MSMs, having simpler SumChecks, and using a 64-bit Goldilocks-64 prime field that yields smaller modmuls. Consequently, NoCap runs all operations several times, including SumChecks 3 times, to obtain 128 bits of security. In contrast, zkSpeed supports arbitrary 255-bit and 381-bit primes for MLEs and elliptic curves points, respectively. We further compare zkSpeed with an iso-area SZKP (Groth16) implementation, giving them the benefit of zkSpeed’s improved MSMs, and optimistically scale up their design to use the BLS12-381 curve. This design, SZKP+, enjoys a 6 reduction in proving time compared to zkSpeed, largely because it has fewer MSMs on its critical path. These speedups come at the cost of circuit-specific setup, incurring large costs any time the application is updated. In sum, NoCap, SZKP, and zkSpeed address different application domains, representing a range of tradeoffs ranging from security properties to software/hardware costs.
Jellyfish: Jellyfish is a HyperPlonk variant supporting gates of arity (fan-in) higher than 2. Unlike R1CS, it supports higher degree constraints, e.g. . The additional expressiveness means, iso-application, the total size of all MLE tables decreases (the number of tables increases with arity, but table size decreases super-proportionally). High-degree gates have utility in many applications (Dellepere et al., 2024); this is especially pronounced when proving the correctness of cryptographic operations like encryption (Xiong et al., 2023) or hash-functions (Grassi et al., 2019). zkSpeed could be extended to support Jellyfish, in which case the ratio of table count to table size may improve the runtime (with sufficient bandwidth). We leave this for future work.
9. Conclusion
This paper presents zkSpeed, the first work to accelerate HyperPlonk proofs in hardware, which offers time complexity compared to prominent zkSNARKs that rely on computational primitives that have complexity (e.g. Groth16). zkSpeed constitutes accelerator units for all core HyperPlonk functions, with special attention paid to prominent kernels: SumCheck and MSM. zkSpeed is a modular architecture, and we leverage a performance model to conduct design space optimization, and analyze the pareto frontier to identify well performing designs. A zkSpeed accelerator with 366 mm2 and 2 TB/s of bandwidth achieves geomean speedup of over CPU baselines, demonstrating the promise zkSpeed offers to accelerate HyperPlonk.
Acknowledgements.
This research was developed with funding in part from the NSF CAREER award #2340137, an NSF RINGS Award, from DARPA under the Data Protection in Virtual Environments (DPRIVE) program, grant number HR0011-21-9-0003, and the Hybrid Electromagnetic side-channel and Interactive-proof Methods to Detect and Amend LogicaL Rifts (HEIMDALLR) programs, grant number HR0011-25-C-0300, and from a gift award from Google.References
- (1)
- lib (2018) 2018. libsnark: a C++ library for zkSNARK proofs. https://github.com/scipr-lab/libsnark
- hbm (2023) 2023. High Bandwidth Memory DRAM (HBM3). Technical Report JESD238A. JEDEC. https://www.jedec.org/standards-documents/docs/jesd238a.
- cer (2023) 2023. What is the ZCash Ceremony? The Complete Beginners Guide. https://coinbureau.com/education/zcash-ceremony/.
- Abbaszadeh et al. (2024) Kasra Abbaszadeh, Christodoulos Pappas, Jonathan Katz, and Dimitrios Papadopoulos. 2024. Zero-Knowledge Proofs of Training for Deep Neural Networks. In Proceedings of the 2024 on ACM SIGSAC Conference on Computer and Communications Security (Salt Lake City, UT, USA) (CCS ’24). Association for Computing Machinery, New York, NY, USA, 4316–4330. https://doi.org/10.1145/3658644.3670316
- Ahmed et al. (2024) Anees Ahmed, Nojan Sheybani, Davi Moreno, Nges Brian Njungle, Tengkai Gong, Michel Kinsy, and Farinaz Koushanfar. 2024. AMAZE: Accelerated MiMC Hardware Architecture for Zero-Knowledge Applications on the Edge. arXiv:2411.06350 [cs.CR] https://arxiv.org/abs/2411.06350 Accepted to ICCAD 2024.
- AMD (2024) Product AMD. 2024. Server Processor Specifications. https://www.amd.com/en/products/specifications/server-processor.html
- Ben-Sasson et al. (2014) Eli Ben-Sasson, Alessandro Chiesa, Christina Garman, Matthew Green, Ian Miers, Eran Tromer, and Madars Virza. 2014. Zerocash: Decentralized Anonymous Payments from Bitcoin. In 2014 IEEE Symposium on Security and Privacy, SP 2014, Berkeley, CA, USA, May 18-21, 2014. IEEE Computer Society, 459–474. https://doi.org/10.1109/SP.2014.36
- Berrut and Trefethen (2004) Jean-Paul Berrut and Lloyd N Trefethen. 2004. Barycentric lagrange interpolation. SIAM review 46, 3 (2004), 501–517.
- Bünz and Chen (2023) Benedikt Bünz and Binyi Chen. 2023. Protostar: Generic Efficient Accumulation/Folding for Special-Sound Protocols. In Advances in Cryptology - ASIACRYPT 2023 - 29th International Conference on the Theory and Application of Cryptology and Information Security, Guangzhou, China, December 4-8, 2023, Proceedings, Part II (Lecture Notes in Computer Science, Vol. 14439), Jian Guo and Ron Steinfeld (Eds.). Springer, 77–110. https://doi.org/10.1007/978-981-99-8724-5_3
- Butt et al. (2024) Shahzad Ahmad Butt, Benjamin Reynolds, Veeraraghavan Ramamurthy, Xiao Xiao, Pohrong Chu, Setareh Sharifian, Sergey Gribok, and Bogdan Pasca. 2024. if-ZKP: Intel FPGA-Based Acceleration of Zero Knowledge Proofs. arXiv:2412.12481 [cs.AR] https://arxiv.org/abs/2412.12481
- Chen et al. (2022) Binyi Chen, Benedikt Bünz, Dan Boneh, and Zhenfei Zhang. 2022. HyperPlonk: Plonk with Linear-Time Prover and High-Degree Custom Gates. Cryptology ePrint Archive, Paper 2022/1355. https://eprint.iacr.org/2022/1355
- Chen et al. (2024) Xiangren Chen, Bohan Yang, Wenping Zhu, Hanning Wang, Qichao Tao, Shuying Yin, Min Zhu, Shaojun Wei, and Leibo Liu. 2024. A High-performance NTT/MSM Accelerator for Zero-knowledge Proof Using Load-balanced Fully-pipelined Montgomery Multiplier. IACR Transactions on Cryptographic Hardware and Embedded Systems 2025, 1 (Dec. 2024), 275–313. https://doi.org/10.46586/tches.v2025.i1.275-313
- Chillotti et al. (2020) Ilaria Chillotti, Nicolas Gama, Mariya Georgieva, and Malika Izabachène. 2020. TFHE: fast fully homomorphic encryption over the torus. Journal of Cryptology 33, 1 (2020), 34–91.
- Daftardar et al. (2024) Alhad Daftardar, Brandon Reagen, and Siddharth Garg. 2024. SZKP: A Scalable Accelerator Architecture for Zero-Knowledge Proofs. In Proceedings of the 2024 International Conference on Parallel Architectures and Compilation Techniques. 271–283.
- de Castro et al. (2021) Leo de Castro, Rashmi Agrawal, Rabia Yazicigil, Anantha Chandrakasan, Vinod Vaikuntanathan, Chiraag Juvekar, and Ajay Joshi. 2021. Does Fully Homomorphic Encryption Need Compute Acceleration? Cryptology ePrint Archive, Paper 2021/1636. https://eprint.iacr.org/2021/1636
- Dellepere et al. (2024) Michel Dellepere, Pratyush Mishra, and Alireza Shirzad. 2024. Garuda and Pari: Faster and Smaller SNARKs via Equifficient Polynomial Commitments. Cryptology ePrint Archive, Paper 2024/1245. https://eprint.iacr.org/2024/1245
- Diamond and Posen (2023) Benjamin E. Diamond and Jim Posen. 2023. Succinct Arguments over Towers of Binary Fields. Cryptology ePrint Archive, Paper 2023/1784. https://eprint.iacr.org/2023/1784
- Diamond and Posen (2024) Benjamin E. Diamond and Jim Posen. 2024. Polylogarithmic Proofs for Multilinears over Binary Towers. IACR Cryptol. ePrint Arch. (2024), 504. https://eprint.iacr.org/2024/504
- Ebel et al. (2025) Austin Ebel, Karthik Garimella, and Brandon Reagen. 2025. Orion: A Fully Homomorphic Encryption Framework for Deep Learning. In Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (Rotterdam, Netherlands) (ASPLOS ’25). Association for Computing Machinery, New York, NY, USA, 734–749. https://doi.org/10.1145/3676641.3716008
- Ebel and Reagen (2024) Austin Ebel and Brandon Reagen. 2024. Osiris: A Systolic Approach to Accelerating Fully Homomorphic Encryption. arXiv:2408.09593 [cs.CR] https://arxiv.org/abs/2408.09593
- Feng et al. (2024) Boyuan Feng, Zheng Wang, Yuke Wang, Shu Yang, and Yufei Ding. 2024. ZENO: A Type-based Optimization Framework for Zero Knowledge Neural Network Inference. In Proceedings of the 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (La Jolla, CA, USA) (ASPLOS ’24). Association for Computing Machinery, New York, NY, USA, 450–464. https://doi.org/10.1145/3617232.3624852
- Gabizon et al. (2019) Ariel Gabizon, Zachary J. Williamson, and Oana Ciobotaru. 2019. PLONK: Permutations over Lagrange-bases for Oecumenical Noninteractive arguments of Knowledge. Cryptology ePrint Archive, Paper 2019/953. https://eprint.iacr.org/2019/953
- Garimella et al. (2023) Karthik Garimella, Zahra Ghodsi, Nandan Kumar Jha, Siddharth Garg, and Brandon Reagen. 2023. Characterizing and Optimizing End-to-End Systems for Private Inference. In Proceedings of the 28th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 3 (Vancouver, BC, Canada) (ASPLOS 2023). Association for Computing Machinery, New York, NY, USA, 89–104. https://doi.org/10.1145/3582016.3582065
- Grassi et al. (2019) Lorenzo Grassi, Dmitry Khovratovich, Christian Rechberger, Arnab Roy, and Markus Schofnegger. 2019. Poseidon: A New Hash Function for Zero-Knowledge Proof Systems. Cryptology ePrint Archive, Paper 2019/458. https://eprint.iacr.org/2019/458
- Groth (2016) Jens Groth. 2016. On the Size of Pairing-based Non-interactive Arguments. Cryptology ePrint Archive, Paper 2016/260. https://eprint.iacr.org/2016/260 https://eprint.iacr.org/2016/260.
- Groth et al. (2018) Jens Groth, Markulf Kohlweiss, Mary Maller, Sarah Meiklejohn, and Ian Miers. 2018. Updatable and Universal Common Reference Strings with Applications to zk-SNARKs. In Advances in Cryptology - CRYPTO 2018 - 38th Annual International Cryptology Conference, Santa Barbara, CA, USA, August 19-23, 2018, Proceedings, Part III (Lecture Notes in Computer Science, Vol. 10993), Hovav Shacham and Alexandra Boldyreva (Eds.). Springer, 698–728. https://doi.org/10.1007/978-3-319-96878-0_24
- Hirner et al. (2025) Florian Hirner, Florian Krieger, and Sujoy Sinha Roy. 2025. Chiplet-Based Techniques for Scalable and Memory-Aware Multi-Scalar Multiplication. Cryptology ePrint Archive, Paper 2025/252. https://eprint.iacr.org/2025/252
- Inc. (2020) Rambus Inc. 2020. White Paper: HBM2E and GDDR6: Memory Solutions for AI. White Paper.
- Jha and Reagen (2024) Nandan Kumar Jha and Brandon Reagen. 2024. DeepReShape: Redesigning Neural Networks for Efficient Private Inference. Transactions on Machine Learning Research (TMLR) (2024).
- Ji et al. (2024) Zhuoran Ji, Zhiyuan Zhang, Jiming Xu, and Lei Ju. 2024. Accelerating Multi-Scalar Multiplication for Efficient Zero Knowledge Proofs with Multi-GPU Systems. In Proceedings of the 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 3 (La Jolla, CA, USA) (ASPLOS ’24). Association for Computing Machinery, New York, NY, USA, 57–70. https://doi.org/10.1145/3620666.3651364
- Kim et al. (2023) Jongmin Kim, Sangpyo Kim, Jaewan Choi, Jaiyoung Park, Donghwan Kim, and Jung Ho Ahn. 2023. SHARP: A Short-Word Hierarchical Accelerator for Robust and Practical Fully Homomorphic Encryption. In Proceedings of the 50th Annual International Symposium on Computer Architecture (Orlando, FL, USA) (ISCA ’23). Association for Computing Machinery, New York, NY, USA, Article 18, 15 pages. https://doi.org/10.1145/3579371.3589053
- Kim et al. (2022b) Jongmin Kim, Gwangho Lee, Sangpyo Kim, Gina Sohn, Minsoo Rhu, John Kim, and Jung Ho Ahn. 2022b. ARK: Fully Homomorphic Encryption Accelerator with Runtime Data Generation and Inter-Operation Key Reuse. In 2022 55th IEEE/ACM International Symposium on Microarchitecture (MICRO). 1237–1254. https://doi.org/10.1109/MICRO56248.2022.00086
- Kim et al. (2022a) Sangpyo Kim, Jongmin Kim, Michael Jaemin Kim, Wonkyung Jung, John Kim, Minsoo Rhu, and Jung Ho Ahn. 2022a. BTS: An Accelerator for Bootstrappable Fully Homomorphic Encryption. In Proceedings of the 49th Annual International Symposium on Computer Architecture (New York, New York) (ISCA ’22). Association for Computing Machinery, New York, NY, USA, 711–725. https://doi.org/10.1145/3470496.3527415
- Kohrita and Towa (2024) Tohru Kohrita and Patrick Towa. 2024. Zeromorph: Zero-Knowledge Multilinear-Evaluation Proofs from Homomorphic Univariate Commitments. J. Cryptol. 37, 4 (2024), 38. https://doi.org/10.1007/S00145-024-09519-0
- Kothapalli et al. (2022) Abhiram Kothapalli, Srinath T. V. Setty, and Ioanna Tzialla. 2022. Nova: Recursive Zero-Knowledge Arguments from Folding Schemes. In Advances in Cryptology - CRYPTO 2022 - 42nd Annual International Cryptology Conference, CRYPTO 2022, Santa Barbara, CA, USA, August 15-18, 2022, Proceedings, Part IV (Lecture Notes in Computer Science, Vol. 13510), Yevgeniy Dodis and Thomas Shrimpton (Eds.). Springer, 359–388. https://doi.org/10.1007/978-3-031-15985-5_13
- Liu et al. (2024a) Changxu Liu, Hao Zhou, Patrick Dai, Li Shang, and Fan Yang. 2024a. PriorMSM: An Efficient Acceleration Architecture for Multi-Scalar Multiplication. ACM Transactions on Design Automation of Electronic Systems 29, 5 (2024), 1–26.
- Liu et al. (2024b) Changxu Liu, Hao Zhou, Lan Yang, Zheng Wu, Patrick Dai, Yinlong Li, Shiyong Wu, and Fan Yang. 2024b. Myosotis: An Efficiently Pipelined and Parameterized Multi-Scalar Multiplication Architecture via Data Sharing. IEEE Transactions on Computer-Aided Design of Integrated Circuits and Systems (2024), 1–1. https://doi.org/10.1109/TCAD.2024.3524364
- Liu et al. (2024c) Changxu Liu, Hao Zhou, Lan Yang, Jiamin Xu, Patrick Dai, and Fan Yang. 2024c. Gypsophila: A Scalable and Bandwidth-Optimized Multi-Scalar Multiplication Architecture. In Proceedings of the 61st ACM/IEEE Design Automation Conference (San Francisco, CA, USA) (DAC ’24). Association for Computing Machinery, New York, NY, USA, Article 94, 6 pages. https://doi.org/10.1145/3649329.3658259
- Lu et al. (2024) Tao Lu, Yuxun Chen, Zonghui Wang, Xiaohang Wang, Wenzhi Chen, and Jiaheng Zhang. 2024. BatchZK: A Fully Pipelined GPU-Accelerated System for Batch Generation of Zero-Knowledge Proofs. Cryptology ePrint Archive, Paper 2024/1862. https://eprint.iacr.org/2024/1862
- Lu et al. (2022) Tao Lu, Chengkun Wei, Ruijing Yu, Chaochao Chen, Wenjing Fang, Lei Wang, Zeke Wang, and Wenzhi Chen. 2022. cuZK: Accelerating Zero-Knowledge Proof with A Faster Parallel Multi-Scalar Multiplication Algorithm on GPUs. Cryptology ePrint Archive, Paper 2022/1321. https://eprint.iacr.org/2022/1321 https://eprint.iacr.org/2022/1321.
- Ma et al. (2023) Weiliang Ma, Qian Xiong, Xuanhua Shi, Xiaosong Ma, Hai Jin, Haozhao Kuang, Mingyu Gao, Ye Zhang, Haichen Shen, and Weifang Hu. 2023. GZKP: A GPU Accelerated Zero-Knowledge Proof System. In Proceedings of the 28th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (Vancouver, BC, Canada) (ASPLOS 2023). Association for Computing Machinery, New York, NY, USA, 340–353. https://doi.org/10.1145/3575693.3575711
- Mo et al. (2025a) Jianqiao Mo, Alhad Daftardar, Joey Ah-kiow, Kaiyue Guo, Benedikt Bünz, Siddharth Garg, and Brandon Reagen. 2025a. MTU: The Multifunction Tree Unit in zkSpeed for Accelerating HyperPlonk. arXiv:2507.16793 [cs.AR] https://arxiv.org/abs/2507.16793
- Mo et al. (2023) Jianqiao Mo, Jayanth Gopinath, and Brandon Reagen. 2023. HAAC: A Hardware-Software Co-Design to Accelerate Garbled Circuits. In Proceedings of the 50th Annual International Symposium on Computer Architecture (Orlando, FL, USA) (ISCA ’23). Association for Computing Machinery, New York, NY, USA, Article 10, 13 pages. https://doi.org/10.1145/3579371.3589045
- Mo et al. (2025b) Jianqiao Cambridge Mo, Karthik Garimella, Austin Ebel, and Brandon Reagen. 2025b. ABLE: Optimizing Mixed Arithmetic and Boolean Garbled Circuit. Cryptology ePrint Archive, Paper 2025/048. https://eprint.iacr.org/2025/048
- Monnot (2024) Barnabé Monnot. 2024. On Block Sizes, Gas Limits, and Scalability. https://ethresear.ch/t/on-block-sizes-gas-limits-and-scalability/18444. Accessed: 2025-05-06.
- Montgomery (1987) Peter L. Montgomery. 1987. Speeding the Pollard and elliptic curve methods of factorization. Math. Comp. 48, 177 (1987), 243–264. https://doi.org/10.1090/S0025-5718-1987-0866113-7
- Morgan (2019a) Timothy Prickett Morgan. 2019a. AMD Doubles Down – And Up – With Rome Epyc Server Chips. https://www.nextplatform.com/2019/08/07/amd-doubles-down-and-up-with-rome-epyc-server-chips/
- Morgan (2019b) Timothy Prickett Morgan. 2019b. A Deep Dive Into AMD’s Rome EPYC Architecture. https://www.nextplatform.com/2019/08/15/a-deep-dive-into-amds-rome-epyc-architecture/
- Neda et al. (2024) Negar Neda, Austin Ebel, Benedict Reynwar, and Brandon Reagen. 2024. CiFlow: Dataflow Analysis and Optimization of Key Switching for Homomorphic Encryption. In 2024 IEEE International Symposium on Performance Analysis of Systems and Software (ISPASS). 61–72. https://doi.org/10.1109/ISPASS61541.2024.00016
- Nelson (2022) Jason Nelson. 2022. Zcash Nixes Trusted Setup, Enters New Era With Major Network Update. https://decrypt.co/101762/zcash-nixes-trusted-setup-enters-new-era-with-major-network-update
- OpenCores (2013) OpenCores. 2013. SHA-3 IP Core. https://opencores.org/projects/sha3. Accessed: November 16, 2024.
- Pippenger (1976) Nicholas Pippenger. 1976. On the evaluation of powers and related problems. In 17th Annual Symposium on Foundations of Computer Science (sfcs 1976). 258–263. https://doi.org/10.1109/SFCS.1976.21
- Pornin (2020) Thomas Pornin. 2020. Optimized Binary GCD for Modular Inversion. 2020/972 (2020). https://eprint.iacr.org/2020/972
- Qiu et al. (2024) Pengcheng Qiu, Guiming Wu, Tingqiang Chu, Changzheng Wei, Runzhou Luo, Ying Yan, Wei Wang, and Hui Zhang. 2024. MSMAC: Accelerating Multi-Scalar Multiplication for Zero-Knowledge Proof. In Proceedings of the 61st ACM/IEEE Design Automation Conference (San Francisco, CA, USA) (DAC ’24). Association for Computing Machinery, New York, NY, USA, Article 66, 6 pages. https://doi.org/10.1145/3649329.3655672
- Samardzic et al. (2021) Nikola Samardzic, Axel Feldmann, Aleksandar Krastev, Srinivas Devadas, Ronald Dreslinski, Christopher Peikert, and Daniel Sanchez. 2021. F1: A Fast and Programmable Accelerator for Fully Homomorphic Encryption. In MICRO-54: 54th Annual IEEE/ACM International Symposium on Microarchitecture (Virtual Event, Greece) (MICRO ’21). Association for Computing Machinery, New York, NY, USA, 238–252. https://doi.org/10.1145/3466752.3480070
- Samardzic et al. (2022) Nikola Samardzic, Axel Feldmann, Aleksandar Krastev, Nathan Manohar, Nicholas Genise, Srinivas Devadas, Karim Eldefrawy, Chris Peikert, and Daniel Sanchez. 2022. CraterLake: A Hardware Accelerator for Efficient Unbounded Computation on Encrypted Data. In Proceedings of the 49th Annual International Symposium on Computer Architecture (New York, New York) (ISCA ’22). Association for Computing Machinery, New York, NY, USA, 173–187. https://doi.org/10.1145/3470496.3527393
- Samardzic et al. (2024) Nikola Samardzic, Simon Langowski, Srinivas Devadas, and Daniel Sanchez. 2024. Accelerating Zero-Knowledge Proofs Through Hardware-Algorithm Co-Design. Preprint. https://people.csail.mit.edu/devadas/pubs/micro24_nocap.pdf.
- Schlachter and Drake (2019) Scott Schlachter and Brian Drake. 2019. Introducing Micron® DDR5 SDRAM: More than a generational update. XP055844818 31 (2019), 6.
- Setty (2020) Srinath Setty. 2020. Spartan: Efficient and General-Purpose zkSNARKs Without Trusted Setup. In Advances in Cryptology – CRYPTO 2020, Daniele Micciancio and Thomas Ristenpart (Eds.). Springer International Publishing, Cham, 704–737.
- Setty et al. (2024) Srinath T. V. Setty, Justin Thaler, and Riad S. Wahby. 2024. Unlocking the Lookup Singularity with Lasso. In Advances in Cryptology - EUROCRYPT 2024 - 43rd Annual International Conference on the Theory and Applications of Cryptographic Techniques, Zurich, Switzerland, May 26-30, 2024, Proceedings, Part VI (Lecture Notes in Computer Science, Vol. 14656), Marc Joye and Gregor Leander (Eds.). Springer, 180–209. https://doi.org/10.1007/978-3-031-58751-1_7
- Sheybani et al. (2025) Nojan Sheybani, Tengkai Gong, Anees Ahmed, Nges Brian Njungle, Michel Kinsy, and Farinaz Koushanfar. 2025. Gotta Hash ’Em All! Speeding Up Hash Functions for Zero-Knowledge Proof Applications. arXiv:2501.18780 [cs.CR] https://arxiv.org/abs/2501.18780
- Soni et al. (2023) Deepraj Soni, Negar Neda, Naifeng Zhang, Benedict Reynwar, Homer Gamil, Benjamin Heyman, Mohammed Nabeel, Ahmad Al Badawi, Yuriy Polyakov, Kellie Canida, Massoud Pedram, Michail Maniatakos, David Bruce Cousins, Franz Franchetti, Matthew French, Andrew Schmidt, and Brandon Reagen. 2023. RPU: The Ring Processing Unit. In 2023 IEEE International Symposium on Performance Analysis of Systems and Software (ISPASS). 272–282. https://doi.org/10.1109/ISPASS57527.2023.00034
- tech (2024) powerup tech. 2024. AMD EPYC 7502 Specs. https://www.techpowerup.com/cpu-specs/epyc-7502.c2250
- Thaler (2013) Justin Thaler. 2013. Time-Optimal Interactive Proofs for Circuit Evaluation. Cryptology ePrint Archive, Paper 2013/351. https://eprint.iacr.org/2013/351
- Thaler (2022) Justin Thaler. 2022. Proofs, Arguments, and Zero-Knowledge. https://people.cs.georgetown.edu/jthaler/ProofsArgsAndZK.pdf
- Valiant (2008) Paul Valiant. 2008. Incrementally Verifiable Computation or Proofs of Knowledge Imply Time/Space Efficiency. In Theory of Cryptography, Fifth Theory of Cryptography Conference, TCC 2008, New York, USA, March 19-21, 2008 (Lecture Notes in Computer Science, Vol. 4948), Ran Canetti (Ed.). Springer, 1–18. https://doi.org/10.1007/978-3-540-78524-8_1
- Wang and Gao (2023) Cheng Wang and Mingyu Gao. 2023. SAM: A Scalable Accelerator for Number Theoretic Transform Using Multi-Dimensional Decomposition. In 2023 IEEE/ACM International Conference on Computer Aided Design (ICCAD). 1–9. https://doi.org/10.1109/ICCAD57390.2023.10323744
- Wang and Gao (2025) Cheng Wang and Mingyu Gao. 2025. UniZK: Accelerating Zero-Knowledge Proof with Unified Hardware and Flexible Kernel Mapping. In Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (Rotterdam, Netherlands) (ASPLOS ’25). Association for Computing Machinery, New York, NY, USA, 1101–1117. https://doi.org/10.1145/3669940.3707228
- Xie et al. (2022) Tiancheng Xie, Yupeng Zhang, and Dawn Song. 2022. Orion: Zero Knowledge Proof with Linear Prover Time. In Advances in Cryptology - CRYPTO 2022 - 42nd Annual International Cryptology Conference, CRYPTO 2022, Santa Barbara, CA, USA, August 15-18, 2022, Proceedings, Part IV (Lecture Notes in Computer Science, Vol. 13510), Yevgeniy Dodis and Thomas Shrimpton (Eds.). Springer, 299–328. https://doi.org/10.1007/978-3-031-15985-5_11
- Xiong et al. (2023) Alex Luoyuan Xiong, Binyi Chen, Zhenfei Zhang, Benedikt Bünz, Ben Fisch, Fernando Krell, and Philippe Camacho. 2023. VeriZexe: Decentralized Private Computation with Universal Setup. In 32nd USENIX Security Symposium, USENIX Security 2023, Anaheim, CA, USA, August 9-11, 2023, Joseph A. Calandrino and Carmela Troncoso (Eds.). USENIX Association, 4445–4462. https://www.usenix.org/conference/usenixsecurity23/presentation/xiong
- Yang et al. (2025) Zhengbang Yang, Lutan Zhao, Peinan Li, Han Liu, Kai Li, Boyan Zhao, Dan Meng, and Rui Hou. 2025. LegoZK: A Dynamically Reconfigurable Accelerator for Zero-Knowledge Proof. In 2025 IEEE International Symposium on High Performance Computer Architecture (HPCA). 113–126. https://doi.org/10.1109/HPCA61900.2025.00020
- Zhang and Franchetti (2025) Naifeng Zhang and Franz Franchetti. 2025. Code Generation for Cryptographic Kernels using Multi-word Modular Arithmetic on GPU. In Proceedings of the 23rd ACM/IEEE International Symposium on Code Generation and Optimization (Las Vegas, NV, USA) (CGO ’25). Association for Computing Machinery, New York, NY, USA, 476–492. https://doi.org/10.1145/3696443.3708948
- Zhang et al. (2021) Ye Zhang, Shuo Wang, Xian Zhang, Jiangbin Dong, Xingzhong Mao, Fan Long, Cong Wang, Dong Zhou, Mingyu Gao, and Guangyu Sun. 2021. PipeZK: Accelerating Zero-Knowledge Proof with a Pipelined Architecture. In 2021 ACM/IEEE 48th Annual International Symposium on Computer Architecture (ISCA).
- Zhang et al. (2022) Zhenfei Zhang, Binyi Chen, Benedikt Bünz, and Alex Xiong. 2022. Hyperplonk Implementation. https://github.com/EspressoSystems/hyperplonk
- Zhou et al. (2024) Hao Zhou, Changxu Liu, Lan Yang, Li Shang, and Fan Yang. 2024. ReZK: A Highly Reconfigurable Accelerator for Zero-Knowledge Proof. IEEE Transactions on Circuits and Systems I: Regular Papers (2024).
- Zhu et al. (2024) Xudong Zhu, Haoqi He, Zhengbang Yang, Yi Deng, Lutan Zhao, and Rui Hou. 2024. Elastic MSM: A Fast, Elastic and Modular Preprocessing Technique for Multi-Scalar Multiplication Algorithm on GPUs. Cryptology ePrint Archive, Paper 2024/057. https://eprint.iacr.org/2024/057